Prove that all the real of the form a+b√2 a,b element of Z forms a ring

1
Expert's answer
2022-01-18T07:35:43-0500

Denote by "K" the set of all real numbers of the form "a+b\sqrt{2}" where "a" and "b" are integers, that is

"K=\{ a+b\sqrt{2}\ |\ a,b\in\mathbb Z\}."


To prove that "(K,+,\cdot)" is a ring it is sufficient to prove that "(K,+,\cdot)" is a subring of the ring "(\mathbb R,+,\cdot)" of real numbers.

Let "a+b\sqrt{2}, c+d\sqrt{2}\in K"

"a+b\sqrt{2}, c+d\sqrt{2}\in K""a+b\sqrt{2}, c+d\sqrt{2}\in K", where "a,b,c,d\in\mathbb Z." Then "(a+b\sqrt{2})+ (c+d\sqrt{2})=(a+c)+(b+d)\sqrt{2}\in K""(a+b\sqrt{2})+ (c+d\sqrt{2})=(a+c)+(b+d)\sqrt{2}\in K""(a+b\sqrt{2})+ (c+d\sqrt{2})=(a+c)+(b+d)\sqrt{2}\in K""(a+b\sqrt{2})+ (c+d\sqrt{2})=(a+c)+(b+d)\sqrt{2}\in K", since "a+c,b+d\in\mathbb Z,"

"(a+b\sqrt{2})- (c+d\sqrt{2})=(a-c)+(b-d)\sqrt{2}\in K"

"(a+b\sqrt{2})- (c+d\sqrt{2})=(a-c)+(b-d)\sqrt{2}\in K""(a+b\sqrt{2})- (c+d\sqrt{2})=(a-c)+(b-d)\sqrt{2}\in K""(a+b\sqrt{2})- (c+d\sqrt{2})=(a-c)+(b-d)\sqrt{2}\in K", since "a-c,b-d\in\mathbb Z,"

"(a+b\sqrt{2})\cdot (c+d\sqrt{2})=(ac+2bd)+(ad+bc)\sqrt{2}\in K"

"(a+b\sqrt{2})\cdot (c+d\sqrt{2})=(ac+2bd)+(ad+bc)\sqrt{2}\in K""(a+b\sqrt{2})\cdot (c+d\sqrt{2})=(ac+2bd)+(ad+bc)\sqrt{2}\in K""(a+b\sqrt{2})\cdot (c+d\sqrt{2})=(ac+2bd)+(ad+bc)\sqrt{2}\in K", since "ac+2bd, ad+bc\in\mathbb Z."

Therefore, "(K,+,\cdot)" is a subring of the ring "(\mathbb R,+,\cdot)", that is "(K,+,\cdot)" a ring.

Since "(\mathbb R,+,\cdot)" is a commutative ring, "(K,+,\cdot)" is a commutative ring as well. The number "1=1+0\sqrt{2}"

"1=1+0\sqrt{2}" is the identity of the ring "(K,+,\cdot)". It follows that "(K,+,\cdot)" is a commutative ring with identity. Since for any real numbers the equality "\alpha\cdot\beta=0" implies that "\alpha=0" or "\beta=0," in particular, for "a+b\sqrt{2}, c+d\sqrt{2}\in K""a+b\sqrt{2}, c+d\sqrt{2}\in K""a+b\sqrt{2}, c+d\sqrt{2}\in K", the equality "(a+b\sqrt{2})(c+d\sqrt{2})=0""(a+b\sqrt{2})(c+d\sqrt{2})=0""(a+b\sqrt{2})(c+d\sqrt{2})=0" implies that "a+b\sqrt{2}=0""a+b\sqrt{2}=0" or "c+d\sqrt{2}=0.""c+d\sqrt{2}=0."

Therefore, "(K,+,\cdot)" is a ring.